OpenStudy (anonymous):

15b+4b-4. need help factoring

OpenStudy (anonymous):

I am assuming you mean 15b^2+4b-4 (5b-2)(3b+2)

jhonyy9 (jhonyy9):

15b+4b-4=15b+4(b-1)=3b+12b+4(b-1)=3b+4(3b+b-1)=3b+4(4b-1)

jhonyy9 (jhonyy9):

or 19b-4